# Invariant Subspace Decomposition

### v1

@@ 3.2 Adaptation task
where $\mathcal{P}$ is the class of distributions over $(X,Y)$ such that $Y=X^{\top}\gamma+\epsilon$, $\mathbb{E}[\epsilon\mid X]=0$ and $\mathbb{E}[\epsilon^{2}\mid X]<\sigma^{2}$. It follows from (17) that, for a generic estimator $\hat{\gamma}_{t^{*}}$ of $\gamma_{0,t^{*}}$ based on the adaptation data alone, we can expect at best to achieve a prediction error of $\sigma_{a}^{2}\frac{p}{m}$, where $\sigma_{a}^{2}$ is the variance of $\epsilon_{t}$ for $t\in\mathcal{I}^{\operatorname{ad}}$ (which is assumed to be constant). We can improve on this if we allow the estimator $\hat{\gamma}_{t^{*}}$ to also depend on the historical data. To see this, observe that we can always decompose $\hat{\gamma}_{t^{*}}=\hat{\beta}+\hat{\delta}_{t^{*}}$ with $\hat{\beta}\in\mathcal{S}^{\operatorname{inv}}$ and $\hat{\delta}_{t^{*}}\in\mathcal{S}^{\operatorname{res}}$. Moreover, under Assumption 2 we can split the expected prediction error of $\hat{\gamma}_{t^{*}}$ at $t^{*}$ accordingly as
@@ 4.1.1 Approximate joint block diagonalization
Given the set of estimated covariance matrices $\{\hat{\Sigma}_{k}\}_{k=1}^{K}$, we now need to estimate an orthogonal transformation $\hat{U}$ that approximately joint block diagonalizes them. We provide an overview of joint block diagonalization methods in Section A.3 in the appendix. In our simulated settings, we solve the approximate joint block diagonalization (AJBD) problem via approximate joint diagonalization (AJD), since we found this approach to represent a good trade off between computational complexity and accuracy. More in detail, similarly to what is proposed by Tichavsky and Koldovsky (2012), we start from the output matrix $V\in\mathbb{R}^{p\times p}$ of the uwedge algorithm by Tichavsky and Yeredor (2008), which solves AJD for the set $\{\hat{\Sigma}_{k}\}_{k=1}^{K}$, i.e., it is such that for all $k\in\{1,\dots,K\}$ the matrix $V^{\top}\hat{\Sigma}_{k}V$ is approximately diagonal. We then use the off-diagonal elements of the approximately diagonalized covariance matrices to identify the common diagonal blocks. As described more in detail in Section A.3, this is achieved by finding an appropriate permutation matrix $P\in\mathbb{R}^{p\times p}$ for the columns of $V$ such that for all $k\in\{1,\dots,K\}$ the matrix $(VP)^{\top}\hat{\Sigma}_{k}(VP)$ is approximately joint block diagonal. The estimated irreducible joint block diagonalizer is then given by $\hat{U}=VP$. We denote by $q^{\hat{U}}_{\max}$ the number of estimated diagonal blocks and by $\{\hat{\mathcal{S}_{j}}\}_{j=1}^{q^{\hat{U}}_{\max}}$ the estimated irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$.
@@ 4.1.2 Estimating the invariant and residual subspaces
We now estimate the invariant and residual subspaces using the estimated irreducible joint block diagonalizer $\hat{U}$. To do so, we first estimate the true time-varying parameter $\gamma_{0,t}$ using similar considerations as the ones made in Section 4.1.1. We assume for example that $\gamma_{0,t}$ is approximately constant in small windows (for simplicity, we consider the same $K$ windows defined in Section 4.1.1). We then compute the regression coefficient $\hat{\gamma}_{k}$ of $\mathbf{Y}_{k}$ on $\mathbf{X}_{k}$, where $\mathbf{Y}_{k}\in\mathbb{R}^{w\times 1}$ and $\mathbf{X}_{k}\in\mathbb{R}^{w\times p}$ are the observations in the $k$-th time window. We have so far omitted the intercept in our linear model. This can always be included by adding a constant term to $X_{t}$ and increasing by one the dimension of $\hat{U}$, adding a row and column at the same index of the constant term, with the off-diagonal entries set to zero and the diagonal entry set to one.
We now use the estimates $\hat{\gamma}_{k}$ to determine which of the subspaces identified by $\hat{U}$ are opt-invariant. For all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$, we take the sets of indices $S_{j}$ as defined in Proposition 1, and consider the estimated orthogonal projection matrices $\Pi_{\hat{\mathcal{S}}_{j}}=\hat{U}^{S_{j}}(\hat{U}^{S_{j}})^{\top}$ from $\mathbb{R}^{p}$ onto the $j$-th subspace $\hat{\mathcal{S}_{j}}$ of the estimated irreducible orthogonal partition. It follows from the proof of Proposition 2 $(i)$ that we can find the opt-invariant subspaces by checking for all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$ whether $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$ remains approximately constant for $k\in\{1,\dots,K\}$.
We further use that, by Lemma 8, if $\hat{\mathcal{S}}_{j}$ is opt-invariant on $[n]$ then the (approximately constant) projected regression coefficient $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$ approximately satisfies the time-invariance constraint (4). This approach is motivated by Proposition 2 $(iii)$. If the corresponding assumption cannot be assumed to hold, other methods can alternatively be used to determine whether $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$ is constant, e.g., checking its gradient or variance. In (4), we can equivalently consider the correlation in place of the covariance, that is, $\operatorname{corr}(Y_{t}-X_{t}^{\top}\beta,X_{t}^{\top}\beta)=0$: an estimate of this correlation allows us to obtain a normalized measure of (4) that is comparable across different experiments. Formally, we consider
$\hat{c}_{k}^{j}\coloneqq\widehat{\operatorname{Corr}}(\mathbf{Y}_{k}-\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}),\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}))$
@@ 4.2 Estimating the invariant and residual components
We consider now the new observation of the covariates $X_{t^{*}}$ at time $t^{*}$ and the adaptation data $(X_{t},Y_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}$ introduced in Section 3, and denote by $\mathbf{X}^{\operatorname{ad}}\in\mathbb{R}^{m\times p}$ and $\mathbf{Y}^{\operatorname{ad}}\in\mathbb{R}^{m\times 1}$ the matrices containing this adaptation data. Similarly to $\hat{\beta}^{\operatorname{inv}}$, using (15) we obtain the following plug-in estimator for $\delta^{\operatorname{res}}_{t^{*}}$
@@ 4.3 Finite sample generalization guarantee
$\operatorname{MSPE}(\gamma^{\operatorname{OLS}}_{t^{*}})-\operatorname{MSPE}(\gamma^{\operatorname{ISD}}_{t^{*}})\geq\sigma_{\operatorname{ad}}^{2}\tfrac{\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}{m}-\sigma_{\epsilon,\max}^{2}\tfrac{\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}{n}C_{\operatorname{inv}}.$
@@
5 Simulation experiments
To show the effectiveness of the ISD framework we report the results of two simulation experiments. The first evaluates the estimation accuracy of the invariant component $\hat{\beta}^{\operatorname{inv}}$ for increasing sample size $n$ of the historical data. The second experiment compares the predictive accuracy of $\hat{\gamma}^{\operatorname{ISD}}_{t}$ and $\hat{\gamma}^{\operatorname{OLS}}_{t}$ for different sizes $m$ of the adaptation dataset, to empirically investigate the dependence of the MSPE difference on the size of the adaptation data shown in Theorem 3.
In both experiments we let the dimension of the covariates be $p=10$, and $\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})=7$, $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})=3$, and generate data as follows. We sample a random orthogonal matrix $U$, and sample the covariates $X_{t}$ from a normal distribution with zero mean and covariance matrix $U\tilde{\Sigma}_{t}U^{\top}$, where $\tilde{\Sigma}_{t}$ is a block-diagonal matrix with four blocks of dimensions $2,4,3$ and $1$, and random entries that change $10$ times in the observed time horizon $n$. We take as true time-varying parameter the rotation by $U$ of the parameter with constant entries equal to $0.2$ (we set these entries all to the same value for simplicity, but they need not to be equal in general) corresponding to the blocks of sizes $4$ and $3$, and time-varying entries—corresponding to the blocks of sizes $2$ and $1$—equal to $(1-1.5t\sin^{2}(it/n+i))/n$, where $i\in\{2,...,8\}$ is the entry index (the values of these coefficients range between $-0.25$ and $1$). The noise terms $\epsilon_{t}$ are sampled i.i.d. from a normal distribution with zero mean and variance $\sigma_{\epsilon_{t}}^{2}=0.64$. The code for the presented experiments is available at https://github.com/mlazzaretto/Invariant-Subspace-Decomposition. The implementation of the uwedge algorithm is taken from the Python package https://github.com/sweichwald/coroICA-python developed by Pfister et al. (2019b).
@@ 5.1 Invariant decomposition and zero-shot prediction
We then consider a separate time window of $250$ observations in which the value of the time-varying coefficients (before the transformation using $U$) is set to $-1$. We use these observations to test the zero-shot predictive capability of the estimated invariant component, i.e., they can be seen as realizations of the variable $X_{t^{*}}$ introduced in Section 3 (with a slight abuse of notation, we also refer to this window as adaptation data even though we do not perform the adaptation step). We compare the predictive performance of the parameter $\hat{\beta}^{\operatorname{inv}}$ on the historical and adaptation data with that of the oracle invariant parameter $\beta^{\operatorname{inv}}$, the maximin effect $\hat{\beta}^{\operatorname{mm}}$ (computed using the magging estimator proposed by  Bühlmann and Meinshausen, 2015), and the OLS solution $\hat{\beta}^{\operatorname{OLS}}$, both computed using the historical data. We show in Figure 5 the results in terms of the $R^{2}$ coefficient, given by $R^{2}=\frac{\sum_{t=1}^{n}(\widehat{\operatorname{Var}}(Y_{t})-\widehat{\operatorname{Var}}(Y_{t}-X_{t}^{\top}\hat{\beta}))}{\sum_{t=1}^{n}\widehat{\operatorname{Var}}(Y_{t})}$.
@@
Figure 5 shows that the $R^{2}$ coefficient of the oracle invariant component $\beta^{\operatorname{inv}}$ remains positive even for values of $\gamma_{0,t}$ in the adaptation data that lie outside of the observed support in the historical data; for increasing $n$ the same holds for the estimated $\hat{\beta}^{\operatorname{inv}}$. Using $\hat{\beta}^{\operatorname{OLS}}$ or $\hat{\beta}^{\operatorname{mm}}$ leads instead to negative explained variance in this experiment.
@@
Figure 5: Normalized explained variance ($R^{2}$) by $\hat{\beta}^{\operatorname{inv}}$ and comparison with $\beta^{\operatorname{inv}}$, $\hat{\beta}^{\operatorname{mm}}$ and $\hat{\beta}^{\operatorname{OLS}}$: training (historical data, left) and zero-shot generalization (test data, right), for different sizes $n$ of the historical data. The dashed line indicates the population value of the (normalized) explained variance by $\beta^{\operatorname{inv}}$.
@@
The main limitation of the ISD method lies in the estimation of the invariant and residual subspaces. As outlined in Section 4, this process consists of two main steps, approximate joint block diagonalization and selection of the invariant blocks, both of which are in practice sensitive to noise. In both steps, we implement our estimator to be as conservative as possible, that is, such that it does not on average overestimate the number of common diagonal blocks or the dimension of $\mathcal{S}^{\operatorname{inv}}$, to avoid including part of the residual subspace into $\hat{\mathcal{S}}^{\operatorname{inv}}$. This behavior is however hard to avoid if the size of the historical dataset is not sufficiently large, therefore requiring large values of $n$ for the ISD framework to work effectively , see Figure 5.

### v2

@@ 3.2 Adaptation task
where $\mathcal{P}$ is the class of distributions over $(X,Y)$ such that $Y=X^{\top}\gamma+\epsilon$, $\mathbb{E}[\epsilon\mid X]=0$ and $\mathbb{E}[\epsilon^{2}\mid X]<\sigma^{2}$. It follows from (18) that, for a generic estimator $\hat{\gamma}_{t^{*}}$ of $\gamma_{0,t^{*}}$ based on the adaptation data alone, we can expect at best to achieve a prediction error of $\sigma_{\operatorname{ad}}^{2}\frac{p}{m}$, where $\sigma_{\operatorname{ad}}^{2}$ is the variance of $\epsilon_{t}$ for $t\in\mathcal{I}^{\operatorname{ad}}$ (which is assumed to be constant). We can improve on this if we allow the estimator $\hat{\gamma}_{t^{*}}$ to also depend on the historical data. To see this, observe that we can always decompose $\hat{\gamma}_{t^{*}}=\hat{\beta}+\hat{\delta}_{t^{*}}$ with $\hat{\beta}\in\mathcal{S}^{\operatorname{inv}}$ and $\hat{\delta}_{t^{*}}\in\mathcal{S}^{\operatorname{res}}$. Moreover, under Assumption 2 we can split the expected prediction error of $\hat{\gamma}_{t^{*}}$ at $t^{*}$ accordingly as
@@ 4.1.1 Approximate joint block diagonalization
Given the set of estimated covariance matrices $\{\hat{\Sigma}_{k}\}_{k=1}^{K}$, we now need to estimate an orthogonal transformation $\hat{U}$ that approximately joint block diagonalizes them. We provide an overview of joint block diagonalization methods in Section A.3 in the appendix. In our simulated settings, we solve the approximate joint block diagonalization (AJBD) problem via approximate joint diagonalization (AJD), since we found this approach to represent a good trade off between computational complexity and accuracy. More in detail, similarly to what is proposed by Tichavsky and Koldovsky (2012), we start from the output matrix $V\in\mathbb{R}^{p\times p}$ of the uwedge algorithm by Tichavsky and Yeredor (2008), which solves AJD for the set $\{\hat{\Sigma}_{k}\}_{k=1}^{K}$, i.e., it is such that for all $k\in\{1,\dots,K\}$ the matrix $V^{\top}\hat{\Sigma}_{k}V$ is approximately diagonal. We then use the off-diagonal elements of the approximately diagonalized covariance matrices to identify the common diagonal blocks. As described more in detail in Section A.3, this is achieved by finding an appropriate permutation matrix $P\in\mathbb{R}^{p\times p}$ for the columns of $V$ such that for all $k\in\{1,\dots,K\}$ the matrix $(VP)^{\top}\hat{\Sigma}_{k}(VP)$ is approximately joint block diagonal. The estimated irreducible joint block diagonalizer is then given by $\hat{U}=VP$. We denote by $q^{\hat{U}}_{\max}$ the number of estimated diagonal blocks and by $\{\hat{\mathcal{S}_{j}}\}_{j=1}^{q^{\hat{U}}_{\max}}$ the estimated irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition.
@@ 4.1.2 Estimating the invariant and residual subspaces
We now estimate the invariant and residual subspaces using the estimated irreducible joint block diagonalizer $\hat{U}$. To do so, we first estimate the true time-varying parameter $\gamma_{0,t}$ using similar considerations as the ones made in Section 4.1.1. We assume for example that $\gamma_{0,t}$ is approximately constant in small windows (for simplicity, we consider the same $K$ windows defined in Section 4.1.1). This assumption is helpful to define the estimation procedure, but is not strictly necessary for the ISD framework to work. We show in Appendix A.2 that the procedure can still work if this assumption is violated. We then compute the regression coefficient $\hat{\gamma}_{k}$ of $\mathbf{Y}_{k}$ on $\mathbf{X}_{k}$, where $\mathbf{Y}_{k}\in\mathbb{R}^{w\times 1}$ and $\mathbf{X}_{k}\in\mathbb{R}^{w\times p}$ are the observations in the $k$-th time window [^4] We have so far omitted the intercept in our linear model, but it can be included by adding a constant term to $X_{t}$ when estimating the linear parameters. We explicitly show how to take the intercept into account in Section B in the appendix.. We use the estimates $\hat{\gamma}_{k}$ to determine which of the subspaces identified by $\hat{U}$ are opt-invariant. For all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$, we take the sets of indices $S_{j}$ as defined in Proposition 1, and consider the estimated orthogonal projection matrices $\Pi_{\hat{\mathcal{S}}_{j}}=\hat{U}^{S_{j}}(\hat{U}^{S_{j}})^{\top}$ from $\mathbb{R}^{p}$ onto the $j$-th subspace $\hat{\mathcal{S}_{j}}$ of the estimated irreducible orthogonal partition. It follows from Lemma 1 that we can find the opt-invariant subspaces by checking for all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$ whether $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$ remains approximately constant for $k\in\{1,\dots,K\}$. To do so, let $\hat{\gamma}\coloneqq(\sum_{k=1}^{K}\operatorname{Var}(\hat{\gamma}_{k})^{-1})^{-1}\sum_{k=1}^{K}\operatorname{Var}(\hat{\gamma}_{k})^{-1}\hat{\gamma}_{k}$ be the average of the estimated regression coefficients inversely weighted by their variance. We further use that, by Lemma 9, if $\hat{\mathcal{S}}_{j}$ is opt-invariant on $[n]$ then the weighted average of the (approximately constant) projected regression coefficient $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$, i.e., $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}$, approximately satisfies the time-invariance constraint (5). This approach is motivated by Proposition 2 $(iii)$. If the corresponding assumption cannot be assumed to hold, other methods can alternatively be used to determine whether $\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}$ is constant, e.g., checking its gradient or variance. In (5), we can equivalently consider the correlation in place of the covariance, that is, $\operatorname{corr}(Y_{t}-X_{t}^{\top}\beta,X_{t}^{\top}\beta)=0$: an estimate of this correlation allows us to obtain a normalized measure of (5) that is comparable across different experiments. Formally, we consider for all $k\in\{1,\dots,K\}$ and for all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$
$\hat{c}_{k}^{j}\coloneqq\widehat{\operatorname{Corr}}(\mathbf{Y}_{k}-\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}),\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}))$
@@
where we approximate opt-invariance with the inequality (19) being satisfied.
@@ 4.2 Estimating the invariant and residual components
We consider now the new observation of the covariates $X_{t^{*}}$ at time $t^{*}$ and the adaptation data $(X_{t},Y_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}$ introduced in Setting 1, and denote by $\mathbf{X}^{\operatorname{ad}}\in\mathbb{R}^{m\times p}$ and $\mathbf{Y}^{\operatorname{ad}}\in\mathbb{R}^{m\times 1}$ the matrices containing this adaptation data. Similarly to $\hat{\beta}^{\operatorname{inv}}$, using (16) we obtain the following plug-in estimator for $\delta^{\operatorname{res}}_{t^{*}}$
@@
We provide the pseudocode summarizing the whole ISD estimation procedure in the Appendix B.
@@ 4.3 Finite sample generalization guarantee
$\operatorname{MSPE}(\hat{\gamma}_{t^{*}}^{\operatorname{OLS}})-\operatorname{MSPE}(\hat{\gamma}_{t^{*}}^{\operatorname{ISD}})\geq\sigma_{\operatorname{ad}}^{2}\tfrac{\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}{m}-\sigma_{\epsilon,\max}^{2}\tfrac{\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}{n}C_{\operatorname{inv}}.$
@@
5 Experiments
To show the effectiveness of the ISD framework we report the results of two simulation experiments and one real data experiment. The first simulation experiment evaluates the estimation accuracy of the invariant component $\hat{\beta}^{\operatorname{inv}}$ for increasing sample size $n$ of the historical data. The second simulation experiment compares the predictive accuracy of $\hat{\gamma}^{\operatorname{ISD}}_{t}$ and $\hat{\gamma}^{\operatorname{OLS}}_{t}$ for different sizes $m$ of the adaptation dataset, to empirically investigate the dependence of the MSPE difference on the size of the adaptation data shown in Theorem 3.
In both simulation experiments we let the dimension of the covariates be $p=10$, and $\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})=7$, $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})=3$, and generate data as follows. We sample a random orthogonal matrix $U$, and sample the covariates $X_{t}$ from a normal distribution with zero mean and covariance matrix $U\tilde{\Sigma}_{t}U^{\top}$, where $\tilde{\Sigma}_{t}$ is a block-diagonal matrix with four blocks of dimensions $2,4,3$ and $1$, and random entries that change $10$ times in the observed time horizon $n$. We take as true time-varying parameter the rotation by $U$ of the parameter with constant entries equal to $0.2$ (we set these entries all to the same value for simplicity, but they need not to be equal in general) corresponding to the blocks of sizes $4$ and $3$, and time-varying entries—corresponding to the blocks of sizes $2$ and $1$—equal to $(1-1.5t\sin^{2}(it/n+i))/n$, where $i\in\{2,...,8\}$ is the entry index (the values of these coefficients range between $-0.25$ and $1$). The noise terms $\epsilon_{t}$ are sampled i.i.d. from a normal distribution with zero mean and variance $\sigma_{\epsilon_{t}}^{2}=0.64$. The dimensions of the subspaces, $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$, and the true time-varying parameter $\gamma_{0,t}$ are chosen to ensure that, in the historical data, $X_{t}^{\top}\beta^{\operatorname{inv}}$ and $X_{t}^{\top}\delta_{t}^{\operatorname{res}}$ explain approximately half of the variance of $Y_{t}$ each. This choice allows for better visualization in the described experiments.
In Section 5.3 we repeat the same experiments on real data coming from a controlled physical system (light tunnel) developed by Gamella et al. (2025). The data used in this experiment will be soon available at https://github.com/juangamella/causal-chamber. The code for the presented experiments is available at https://github.com/mlazzaretto/Invariant-Subspace-Decomposition. The implementation of the uwedge algorithm is taken from the Python package https://github.com/sweichwald/coroICA-python developed by Pfister et al. (2019b).
@@ 5.1 Invariant decomposition and zero-shot prediction
We then consider a separate time window of $250$ observations in which the value of the time-varying coefficients (before the transformation using $U$) is set to $-1$.
Figure 4: MSE of $\hat{\beta}^{\operatorname{inv}}$ for increasing size of the historical data $n$ (see Section 5.1). For larger values of $n$, the estimation of the invariant subspace decomposition becomes more precise and leads to smaller errors in the estimated invariant component $\hat{\beta}^{\operatorname{inv}}$.
@@
We use these observations to test the zero-shot predictive capability of the estimated invariant component, i.e., they can be seen as realizations of the variable $X_{t^{*}}$ introduced in Setting 1 (we refer to this window as test data). We compare the predictive performance of the parameter $\hat{\beta}^{\operatorname{inv}}$ on the historical and test data with that of the oracle invariant parameter $\beta^{\operatorname{inv}}$, the maximin effect $\hat{\beta}^{\operatorname{mm}}$ (computed using the magging estimator proposed by  Bühlmann and Meinshausen, 2015), and the OLS solution $\hat{\beta}^{\operatorname{OLS}}$, both computed using the historical data. We show in Figure 5 the results in terms of the $R^{2}$ coefficient, given by $R^{2}=\frac{\sum_{t=1}^{n}(\widehat{\operatorname{Var}}(Y_{t})-\widehat{\operatorname{Var}}(Y_{t}-X_{t}^{\top}\hat{\beta}))}{\sum_{t=1}^{n}\widehat{\operatorname{Var}}(Y_{t})}$.
@@
Figure 5 shows that the $R^{2}$ coefficient of the oracle invariant component $\beta^{\operatorname{inv}}$ remains positive even for values of $\gamma_{0,t}$ in the test data that lie outside of the observed support in the historical data; for increasing $n$ the same holds for the estimated $\hat{\beta}^{\operatorname{inv}}$. Using $\hat{\beta}^{\operatorname{OLS}}$ or $\hat{\beta}^{\operatorname{mm}}$ leads instead to negative explained variance in this experiment.
@@
Figure 5: Normalized explained variance ($R^{2}$) by $\hat{\beta}^{\operatorname{inv}}$ and comparison with $\beta^{\operatorname{inv}}$, $\hat{\beta}^{\operatorname{mm}}$ and $\hat{\beta}^{\operatorname{OLS}}$: training (historical data, left) and zero-shot generalization (test data, right), for different sizes $n$ of the historical data (see Section 5.1). The dashed line indicates the population value of the (normalized) explained variance by $\beta^{\operatorname{inv}}$.
@@
The main limitation of the ISD method lies in the estimation of the invariant and residual subspaces. As outlined in Section 4, this process consists of two main steps, approximate joint block diagonalization and selection of the invariant blocks, both of which are in practice sensitive to noise. In both steps, we implement our estimator to be as conservative as possible, that is, such that it does not on average overestimate the number of common diagonal blocks or the dimension of $\mathcal{S}^{\operatorname{inv}}$, to avoid including part of the residual subspace into $\hat{\mathcal{S}}^{\operatorname{inv}}$. This behavior is however hard to avoid if the size of the historical dataset is not sufficiently large, therefore requiring large values of $n$ for the ISD framework to work effectively, see Figures 4 and 5.